Saturday, August 4, 2012

Night Log 2012-08-03

SA: Paul
SO: Thea
Others: Charl


* cirrus for the whole night. seeing was moderate for most of the night, 1.4'' -- 1.8''
* had to reject some observations due to observational requirements not met.

* obtained data for the following programs:
- 2012-1-DC_RSA_OTH-001
- 2012-1-POL-001 x2
- 2012-1-UKSC-007
- 2012-1-UNC-001 (very faint source, not sure if conditions were met)
- 2012-1-POL-007
- 2012-1-RSA-019
- 2012-1-RSA_UKSC-006

* spent some time trying to get FP to work, but wasn't successful. I suspect that the etalon needed more time to settle.

No comments:

Post a Comment